Calculate the resistance of a piece of wire of length 200cm given that it's resistivity is 5.4x10^-9 ohm meter. If the cross sectional area of the wire is 9.5x10^-5 cm^2


1
Expert's answer
2021-07-05T08:45:06-0400

Given:

"l=2.0\:\rm m"

"\rho=5.4*10^{-9}\:\rm \Omega\cdot m"

"A=9.5*10^{-9}\:\rm m^2"

The resistance of a wire

"R=\rho\frac{l}{A}"

"=5.4*10^{-9}\:\rm \Omega\cdot m*\frac{2.0\:\rm m}{9.5*10^{-9}\:\rm m^2}=1.1\:\rm \Omega"